OpenStudy (anonymous):

If a = -12/5, then 5x + 2y = 6 and 3x - ay = 4 are parallel. True or False???

OpenStudy (anonymous):

@fiercesteak @Topdog

OpenStudy (anonymous):

5x + 2y = 6 can be rewritten as y = -5x/2 + 3. 3x + 12y/5 =4 becomes y = 5/3 - 15x/12. Therefore 15/12 and -5/2 are different slopes. Which is not parallel

OpenStudy (anonymous):

False. Thanks.. :) @mercer44